Project

General

Profile

NDOS Commissioning Meeting 2011/03/17

Attendance: Jon, Mark Mathis, Tricia, Mark Messier, Enrique, Gavin, Minerba, Leon, Jarek, Denis, Zelimir

Agenda:

  1. DSO Scan Studies
  2. NDOS Installation Update
  3. DAQ/Electronics Plans
  4. DCS Update
  5. AOB

DSO Scans:

  • Mark updated script to use threshold calculation instead of RMS
  • Plots are being updated continuously
  • Working on printing out list of channels that seem to flip frequently
  • Denis: what about long DSO scan? Mark: 14 were taken back-to-back, plots were generated last week, but have not been updated with new threshold calculations

NDOS Installation/Filling:

  • Minerba: could not start filling yesterday as the oil had to be qualified first.
  • Filling began this afternoon.

DAQ/Electronics:

  • Leon:
    • still has not heard if tool for installing noise abatement boards has been delivered to IU
    • DCMs have been installed on 3rd diblock, not yet checked out
  • Jon:
    • DAQ software: investigations into DCM "wig-outs" are underway

DCS

AOB